OpenStudy (anonymous):

a farmer has 2400 feet of fence for a ractangular field along a straght river. no fence is neeeded along the reiver. using calculus find the dimensions of the field which has the largest area

Directrix (directrix):

Let x be the width of the rectangular field. The perimeter of 3 sides of the field (river side not included) = 2400 - x -x = 2400 - 2x. Area = x (2400 - 2x) A(x) = 2400x - 2 x^2 A' (x) = 2400 - 4x 2400 - 4x = 0 x = 600 Dimensions: 600 by 1200

OpenStudy (anonymous):

is the answer 600 * 1200

Directrix (directrix):

The question is to "find the dimensions." In the case of a rectangle, that would be one dimension by the other. It would not the the product which the * symbol indicates. Let's leave it at width: 600 and length: 1200